If x = –0.5, which is the least of these five numbers?
A.x B.x^2 C.x^3 D.x^4 E.x^5

Answer by CharlesG2(834)

If x = –0.5, which is the least of these five numbers?
A.x B.x^2 C.x^3 D.x^4 E.x^5
x = -0.5 = -1/2
x^2 = 1/4
x^3 = -1/8
x^4 = 1/16
x^5 = -1/32 (smallest)
answer is E which is x^5